OpenStudy (anonymous):

translate the equation.... f $0.30 folders and n $1.50 notebooks total $12 grapg the equation and use the graph to determine 3 different combinations of folders an notebooks that total $12

OpenStudy (anonymous):

0.3x + 1.5y = 12

OpenStudy (anonymous):

2*8

OpenStudy (anonymous):

3f + 15n = 120 f + 5n = 40 ->f = - 5n + 40 Can you graph it now?

OpenStudy (anonymous):

i think i can i just dont get this kind of math thats all

OpenStudy (anonymous):

It's linear equation: y = mx + b

OpenStudy (anonymous):

ok isee

OpenStudy (anonymous):

Does it mean you can graph it now?

OpenStudy (anonymous):

im going to try